solve by any method you prefer.
x^2 + 8x + 1 = 0

Answer by fractalier(2101)

Let complete the square on this one...thus from
x^2 + 8x + 1 = 0 we add 15 to both sides
x^2 + 8x + 16 = 15
(x + 4)^2 = 15
x + 4 = ± √15
x = -4 ± √15